# When Attention Sink Emerges in Language Models: An Empirical View

### v1

@@ 3.4 Attention sink under different LMs
Block-wise and head-wise property. We visualize distributions of attention sink across different blocks and heads in LLaMA2/LLaMA3/GPT2/Mistral/Pythia/OPT family in Appendix B.2. We find that (i) though different LMs may have different distributions, they tend to have less attention sink heads in the earlier blocks; (ii) instruction tuning does not significantly modify such distribution.
@@
We pre-train a series of LLaMA models to conduct our experiments. Due to the intractability of replicating LLaMA pre-training, we design small-sized models. Following Liu et al. (2024a), we set hidden dimension $d=\textrm{768}$, block number $L=\textrm{10}$, head number $H=\textrm{8}$, intermediate size of FFN as 1536, resulting in approximately 60M parameters except for word embeddings and unembeddings. We keep the other design the same as LLaMA2 models, including Rotary (Su et al., 2024), pre-norm structure, RMSNorm (Zhang & Sennrich, 2019) as LN, SwiGLU activation (Shazeer, 2020) in FFN, etc.
@@
Learning rate. With a smaller learning rate, it takes longer training steps to lower training loss, as present in Figure 4(Right). Meanwhile, the emergence of attention sink is also delayed. Besides, we also find that a smaller learning rate (1e-4) results in LMs with less obvious attention sink. But further decreasing learning rate significantly affects the optimization and model performance, thus affecting the emergence of attention sink.
@@ 5 Effects of data distribution $p_{\textrm{data}}$ on attention sink
Training data amount. In the default setup, we consider 5B tokens. We wonder whether the attention sink emerges if we further constrain the data within a fixed compute budget. Therefore, we constrain the training data to 5B, 2.5B, 500M, 200M, 100M, and 50M. Meanwhile, we fix the batch size and optimization steps. As visualized in Figure 5(Right), with less training data, the trained LLMs tend to overfit. Meanwhile, attention sink also disappears.
@@
Fixing token in a specific position. Xiao et al. (2023) considered a learnable token in the first token position within each chunk, which can be considered as $\bm{x}_{1}\sim\mathbb{I}(\bm{x}=\bm{x}_{\textrm{fix}})$. We also consider fixing the token $\bm{x}_{\textrm{fix}}$ in the second/third token position during pre-training. Consequently, the attention sink always appears in the fixed token instead of the first token, as shown in Table 8(Right).
@@ 7 Effects of model architecture $p_{\theta}$ on attention sink
In this section, we mainly explore the effects of positional embedding, pre-norm or post-norm structure, and attention design on the emergence of attention sink. In Appendix C, we also show that varying activation functions in the FFN, multi-head design (including concatenation operation and number of heads) do not affect the emergence of attention sink.
@@ 7.1 Positional embedding
Attention sink always appears on the first token, which motivates us to explore where such a position property is brought by positional embedding (PE). Therefore, we attempt to replace the original Rotary with other PEs, as shown in Table 3. We differentiate these PEs through the calculations of the hidden states $\bm{H}^{0}$ before the first transformer block and the dot product between queries and keys $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle$. The detailed formulations are delayed to Appendix A. From the same Table, we observe that only the model with relative PE is difficult to train while other models have comparable performance under our setup. Then we note that all these LMs, even the one without explicit PE (NoPE), have attention sink.
@@ 7.2 Pre-norm and post-norm structure
NoPE $\bm{X}\bm{W}_{E}$ $\bm{q}_{i}\bm{k}_{j}^{T}$ 20.35 3.81
Absolute PE $\bm{X}\bm{W}_{E}+\bm{P}_{\textrm{abs}}$ $\bm{q}_{i}\bm{k}_{j}^{T}$ 32.73 3.74
Learnable PE $\bm{X}\bm{W}_{E}+\bm{P}_{\textrm{learnable}}$ $\bm{q}_{i}\bm{k}_{j}^{T}$ 33.13 3.79
Relative PE $\bm{X}\bm{W}_{E}$ $\bm{q}_{i}\bm{k}_{j}^{T}+g_{\textrm{relative}}(i-j)$ 35.53 5.45
ALiBi $\bm{X}\bm{W}_{E}$ $\bm{q}_{i}\bm{k}_{j}^{T}+g_{\textrm{alibi}}(i-j)$ 20.78 3.71
Rotary $\bm{X}\bm{W}_{E}$ $\bm{q}_{i}\bm{R}_{\Theta\textrm{,}\,j-i}\bm{k}_{j}^{T}$ 18.18 3.73
@@
$\textrm{Softmax}\left(\frac{1}{\sqrt{d_{h}}}\bm{Q}^{l\textrm{,}h}{\bm{K}^{l\textrm{,}h}}^{T}{\sqrt{d_{h}}}+\bm{M}\right)\bm{V}^{l\textrm{,}h}$ 18.18 3.73
$\textrm{Softmax}\left(\frac{1}{\sqrt{d_{h}}}\begin{bmatrix}\bm{q}^{*l,h}\\ \bm{Q}^{l,h}\end{bmatrix}\begin{bmatrix}{\bm{k}^{*l\textrm{,}h}}^{T}&{\bm{K}^{l\textrm{,}h}}^{T}\end{bmatrix}+\bm{M}\right)\begin{bmatrix}\bm{v}^{*l\textrm{,}h}\\ \bm{V}^{l\textrm{,}h}\end{bmatrix}$ 74.12    0.00 3.72
$\textrm{Softmax}\left(\frac{1}{\sqrt{d_{h}}}\bm{Q}^{l\textrm{,}h}\begin{bmatrix}{\bm{k}^{*l\textrm{,}h}}^{T}&{\bm{K}^{l\textrm{,}h}}^{T}\end{bmatrix}+\bm{M}\right)\begin{bmatrix}\bm{v}^{*l\textrm{,}h}\\ \bm{V}^{l\textrm{,}h}\end{bmatrix}$ 72.76    0.04 3.72
$\textrm{Softmax}\left(\frac{1}{\sqrt{d_{h}}}\bm{Q}^{l\textrm{,}h}\begin{bmatrix}{\bm{k}^{*l\textrm{,}h}}^{T}&{\bm{K}^{l\textrm{,}h}}^{T}\end{bmatrix}{\sqrt{d_{h}}}+\bm{M}\right)\begin{bmatrix}\bm{0}\\ \bm{V}^{l\textrm{,}h}\end{bmatrix}$ 73.34    0.00 3.72
$\textrm{Softmax}\left(\frac{1}{\sqrt{d_{h}}}\bm{Q}^{l\textrm{,}h}{\bm{K}^{l\textrm{,}h}}^{T}+\bm{M}\right)\bm{V}^{l\textrm{,}h}+\bm{v}^{*l\textrm{,}h}$ 17.53 3.73
@@ 7.3 Attention biases
Implicit biases in attention. In Section 3.1, we have shown that the first token acts as a bias: its key $\bm{k}_{1}^{l\textrm{,}\,h}$ is distributed in a different manifold and its value $\bm{v}_{1}^{l\textrm{,}\,h}$ has small $\ell_{2}$ norm. Xiao et al. (2023) considered a learnable sink token in each chunk before the input tokens during LM pre-training. As this token is fixed in the first token, this could be considered as implicitly introducing biases $\bm{k}^{*l,h}$, $\bm{v}^{*l,h}$, $\bm{q}^{*l,h}$ in attention, as shown in the second row in Table 4. These biases are the MLP output of $\bm{x}^{*}\bm{W}_{E}$. As a result, attention sink appears on this learnable token $\bm{x}^{*}$ instead of the actual first token $\bm{x}_{1}$.
Explicit biases in attention. Sun et al. (2024) directly introducing $\bm{k}^{*l,h}$ and $\bm{v}^{*l,h}$ as learnable parameters in attention (KV biases). They found that this design could alleviate the massive activations. Considering the important role of $\bm{k}^{*l,h}$ and small $\ell_{2}$-norm of $\bm{v}^{*l,h}$ , we propose introducing only the key biases $\bm{k}^{*l,h}$ and fix value biases $\bm{v}^{*l,h}=\bm{0}$ (K biases). As a control, we also consider only adding value biases (V biases). The formulations of all these attention designs are shown in Table 4.
LMs need key biases. After evaluating the LMs with setups in Table 4, we first observe that these LMs have comparable model performance. Moreover, as long as there are key biases $\bm{k}^{*l\textrm{,}\,h}$, attention sink disappears on the first token but on the biases. From the setup of K biases, we reaffirm that the sink token acts as key biases, storing extra attention scores, which could be completely non-informative and not contribute to the value computation. This is different from the registers (Darcet et al., 2023), which aggregates global information, in Vision Transformers (ViTs) (Dosovitskiy, 2020). It is worth mentioning that the introduced learnable sink token, KV biases, and V biases become part of model parameters in LMs. Removing them will lead to no attention sink in the first position, but a significant drop in model performance. In Figure 7(Middle), we also find that the LM with a learnable sink token has massive activations in $\bm{x}^{*}$. While LMs with KV biases and K biases have no massive activations.
Table 5: (Top) Larger $\ell_{2}$-norm of fixed $\bm{v}^{*l\textrm{,}h}$ results in LMs allocating more attention on $\bm{x}_{1}$ instead of $\bm{k}^{*l\textrm{,}h}$. (Bottom) Even with very few learnable dimensions for $\bm{k}^{*l\textrm{,}h}$, large attention appears in $\bm{k}^{*l\textrm{,}h}$.
@@
$d_{a}$ 1 2 4 8 16 32 64
$\textrm{Sink}_{1}^{\epsilon}$(%)    4.74    4.96    4.39    4.54    2.19    1.94    0.04
valid loss    3.73    3.72    3.72    3.73    3.73    3.73    3.72
Beyond all zeros in V biases. In the setup of K biases, we fix $\bm{v}^{*l,h}=\bm{0}$. We wonder whether the fixed values of $\bm{v}^{*l,h}$ could affect the attention sink. We consider, $\bm{v}^{*l,h}=m\bm{v}^{\prime}$ or $\bm{v}^{*l,h}=m\bm{v}^{\prime\prime}$, where $m$ is the controllable $\ell_{2}$ norm and $\bm{v}^{\prime}=[1,0,0,..,0]$ and $\bm{v}^{\prime\prime}=[1,1,1,..,1]/\sqrt{d_{h}}$. As shown in Table 5(Top), with larger $\ell_{2}$-norm of $\bm{v}^{*l,h}$, attention sink shifts from $\bm{k}^{*l,h}$ to the first token. Intuitively, it is difficult for LMs to remove the effects of $\bm{v}^{*l,h}$ with larger $\ell_{2}$-norm in model predictions. Then they opt to optimize the keys and values of the first token to save extra attention.
Learnable dimensions of K biases. In the setup of K biases, we set all dimensions of $\bm{k}^{*l\textrm{,}\,h}$ as learnable weights. In Section 3.1, we have shown that K biases are distributed in a different manifold, with low rank. Therefore, we consider only $d_{a}$ dimensions of $\bm{k}^{*l\textrm{,}\,h}$ are adjustable/learnable while the other $d_{h}-d_{a}$ dimensions are zeros. As present in Table 5(Bottom), with very few learnable dimensions, even for $d_{a}=\textrm{1}$, $\bm{k}^{*l\textrm{,}\,h}$ are still allocated significant attention. With more learnable dimensions, the attention sink appears more obvious.
Head-sharing biases.. In Appendix C, we explore whether different heads in each block can share the same parameters for KV biases. Results in Table 10(Right) show that the LM with KV biases tends to shift the attention sink from $\bm{k}^{*l,h}$ back to the first token. While the LM with K biases is less affected.
@@ 7.4 Attention operation
General formulation of attention. In the last section, we realize that LMs with softmax attention need key biases to save extra attention. This motivates us to explore whether such a property is related to the dependence among attention scores due to the softmax operation. First, the attention output for the $i$-th token can be generalized into the following formulation: $\bm{v}_{i}^{\dagger}=\left(\sum_{j^{\prime}=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j^{\prime}}))\right)^{-1}\sum_{j=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j}))\bm{v}_{j}=\bm{Z}_{i}^{-1}\sum_{j=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j}))\bm{v}_{j}$, where we omit the positional embedding, and block/head indexes for simplicity. $\bm{Z}_{i}$ is a normalization term and $\varphi(\cdot)$ is a kernel function. Normally, $\bm{Z}_{i}=\sum_{j^{\prime}=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j}))$. For softmax attention, the kernel $\varphi(\cdot)$ is an identity kernel and the similarity function is $\textrm{exp}(\bm{q}_{i}^{T}\bm{k}_{j}/\sqrt{d_{h}})$.
Normalization. In Appendix C, we first show that only scaling the normalization term $\bm{Z}_{i}\rightarrow\bm{Z}_{i}/\alpha$ with $\alpha<\textrm{1}$ results in less obvious attention sink but does not stop its emergence. So we consider removing the normalization. Since the exponential function in softmax tends to explode without normalization, we replace it with sigmoid or elu plus one. When evaluating the attention sink, we compute the proxy attention scores by using the term $\bm{Z}_{i}=\sum_{j^{\prime}=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j^{\prime}}))$ for attention sink metric $\textrm{Sink}_{1}^{\epsilon}$. As shown in Table 6, without normalization, LMs still have comparable validation loss but no attention sink. With normalization, attention sink also emerges in LMs with sigmoid attention.
@@
$\textrm{exp}(\frac{\bm{q}_{i}^{T}\bm{k}_{j}}{\sqrt{d_{h}}})$ $\sum_{j^{\prime}=1}^{i}\textrm{exp}(\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{T}}{\sqrt{d_{h}}})$ 18.18 3.73
$\textrm{sigmoid}(\frac{\bm{q}_{i}\bm{k}_{j}^{T}}{\sqrt{d_{h}}})$ $1$       0.44 3.70
$\textrm{sigmoid}(\frac{\bm{q}_{i}\bm{k}_{j}^{T}}{\sqrt{d_{h}}})$ $\sum_{j^{\prime}=1}^{i}\textrm{sigmoid}(\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{T}}{\sqrt{d_{h}}})$ 30.24 3.74
$\textrm{elu}(\frac{\bm{q}_{i}\bm{k}_{j}^{T}}{\sqrt{d_{h}}})+1$ $1$       0.80 3.69
$\frac{(\textrm{elu}(\bm{q}_{i})+1)(\textrm{elu}(\bm{k}_{j})+1)^{T}}{\sqrt{d_{h}}}$ $\sum_{j^{\prime}=1}^{i}\frac{(\textrm{elu}(\bm{q}_{i})+1)(\textrm{elu}(\bm{k}_{j^{\prime}})+1)^{T}}{\sqrt{d_{h}}}$    53.65 4.19
$\frac{\bm{q}_{i}\bm{k}_{j}^{T}}{\sqrt{d_{h}}}$ $1$      0.00 3.99
$\frac{\textrm{mlp}(\bm{q}_{i})\textrm{mlp}(\bm{k}_{j})^{T}}{\sqrt{d_{h}}}$ $\textrm{max}\left(\left|\sum_{j^{\prime}=1}^{i}\frac{\textrm{mlp}(\bm{q}_{i})\textrm{mlp}(\bm{k}_{j^{\prime}})^{T}}{\sqrt{d_{h}}}\right|\textrm{,}\,1\right)$      0.19 3.85
$\frac{\textrm{mlp}(\bm{q}_{i})\textrm{mlp}(\bm{k}_{j})^{T}}{\sqrt{d_{h}}}$ $1$      0.74 3.91
@@
Scale up to 1B parameters. We compare model behaviors of 1B LMs with softmax attention and sigmoid attention (without normalization). Specifically, the latter achieves a validation loss of 3.10, slightly larger than the 3.07 achieved by the former. However, the attention sink metric significantly drops from 45.11% to near zero: $\textrm{Sink}_{1}^{\epsilon}=\textrm{2.46\%}$% using the proxy attention scores. Meanwhile, as visualized in Figure 8(Right), LLMs with sigmoid attention have no massive activations.

### v2

@@ 3.4 Attention sink under different LMs
We pre-train a series of LLaMA models to conduct our experiments, based on the repos (Zhang et al., 2024a; Liu et al., 2024a). Due to the intractability of replicating LLaMA pre-training, we design small-sized models. Following Liu et al. (2024a), we set hidden dimension $d=\textrm{768}$, block number $L=\textrm{10}$, head number $H=\textrm{8}$, intermediate size of FFN as 1536, resulting in approximately 60M parameters except for word embeddings and unembeddings. We keep the other design the same as LLaMA2 models, including Rotary (Su et al., 2024), pre-norm structure, RMSNorm (Zhang & Sennrich, 2019) as LN, SwiGLU activation (Shazeer, 2020) in FFN, etc.
@@
Learning rate. With a smaller learning rate, it takes longer training steps to lower training loss, as present in Figure 4(Right). Meanwhile, the emergence of attention sink is also delayed. Besides, as shown in Table 9, we also find that a smaller learning rate results in LMs with less obvious attention sink, even if we compensate for more training steps. But further decreasing learning rate significantly affects the optimization and model performance, thus affecting the emergence of attention sink.
@@ 5 Effects of data distribution $p_{\textrm{data}}$ on attention sink
Training data amount. In the default setup, we consider 5B tokens. We wonder whether the attention sink emerges if we further constrain the data within a fixed compute budget. Therefore, we constrain the training data to 5B, 2.5B, 500M, 200M, 100M, and 50M. Meanwhile, we fix the batch size and optimization steps. As visualized in Figure 5(Right), with less training data, attention sink disappears. Further evidence in Figure 28 shows that this is not related to overfitting.
@@
Fixing token in a specific position. Xiao et al. (2023b) considered a learnable token in the first token position within each chunk, which can be considered as $\bm{x}_{1}\sim\mathbb{I}(\bm{x}=\bm{x}_{\textrm{fix}})$. We also consider fixing the token $\bm{x}_{\textrm{fix}}$ in the second/third token position during pre-training. Consequently, the attention sink always appears in the fixed token instead of the first token, as shown in Table 10(Right).
@@ 7 Effects of model architecture $p_{\theta}$ on attention sink
In this section, we mainly explore the effects of positional embedding, pre-norm or post-norm structure, and attention design on the emergence of attention sink. In Appendix D, we also show that varying activation functions in the FFN, multi-head design do not affect the emergence of attention sink.
@@ 7.1 Positional embedding
Attention sink always appears on the first token, which motivates us to explore where such a position property is brought by positional embedding (PE). Therefore, we attempt to replace the original Rotary with other PEs, as shown in Table 3. We differentiate these PEs through the calculations of the hidden states $\bm{H}^{0}$ before the first transformer block and the dot product between queries and keys $\langle\bm{q}_{i}\textrm{,}\,\bm{k}_{j}\rangle$. The detailed formulations are delayed to Appendix B. From the same Table, we observe that only the model with relative PE is difficult to train while other models have comparable performance under our setup. Then we note that all these LMs, even the one without explicit PE (NoPE), have attention sink.
@@ 7.2 Pre-norm and post-norm structure
NoPE $\bm{X}\bm{W}_{E}$ $\bm{q}_{i}\bm{k}_{j}^{\top}$ 20.35 3.81
Absolute PE $\bm{X}\bm{W}_{E}+\bm{P}_{\textrm{abs}}$ $\bm{q}_{i}\bm{k}_{j}^{\top}$ 32.73 3.74
Learnable PE $\bm{X}\bm{W}_{E}+\bm{P}_{\textrm{learnable}}$ $\bm{q}_{i}\bm{k}_{j}^{\top}$ 33.13 3.79
Relative PE $\bm{X}\bm{W}_{E}$ $\bm{q}_{i}\bm{k}_{j}^{\top}+g_{\textrm{relative}}(i-j)$ 35.53 5.45
ALiBi $\bm{X}\bm{W}_{E}$ $\bm{q}_{i}\bm{k}_{j}^{\top}+g_{\textrm{alibi}}(i-j)$ 20.78 3.71
Rotary $\bm{X}\bm{W}_{E}$ $\bm{q}_{i}\bm{R}_{\Theta\textrm{,}\,j-i}\bm{k}_{j}^{\top}$ 18.18 3.73
@@
$\textrm{Softmax}\left(\frac{1}{\sqrt{d_{h}}}\bm{Q}^{l\textrm{,}h}{\bm{K}^{l\textrm{,}h}}^{\top}+\bm{M}\right)\bm{V}^{l\textrm{,}h}$ 18.18 3.73
$\textrm{Softmax}\left(\frac{1}{\sqrt{d_{h}}}\begin{bmatrix}\bm{q}^{*l,h}\\ \bm{Q}^{l,h}\end{bmatrix}\begin{bmatrix}{\bm{k}^{*l\textrm{,}h}}^{\top}&{\bm{K}^{l\textrm{,}h}}^{\top}\end{bmatrix}+\bm{M}\right)\begin{bmatrix}\bm{v}^{*l\textrm{,}h}\\ \bm{V}^{l\textrm{,}h}\end{bmatrix}$ 74.12    0.00 3.72
$\textrm{Softmax}\left(\frac{1}{\sqrt{d_{h}}}\bm{Q}^{l\textrm{,}h}\begin{bmatrix}{\bm{k}^{*l\textrm{,}h}}^{\top}&{\bm{K}^{l\textrm{,}h}}^{\top}\end{bmatrix}+\bm{M}\right)\begin{bmatrix}\bm{v}^{*l\textrm{,}h}\\ \bm{V}^{l\textrm{,}h}\end{bmatrix}$ 72.76    0.04 3.72
$\textrm{Softmax}\left(\frac{1}{\sqrt{d_{h}}}\bm{Q}^{l\textrm{,}h}\begin{bmatrix}{\bm{k}^{*l\textrm{,}h}}^{\top}&{\bm{K}^{l\textrm{,}h}}^{\top}\end{bmatrix}+\bm{M}\right)\begin{bmatrix}\bm{0}\\ \bm{V}^{l\textrm{,}h}\end{bmatrix}$ 73.34    0.00 3.72
$\textrm{Softmax}\left(\frac{1}{\sqrt{d_{h}}}\bm{Q}^{l\textrm{,}h}{\bm{K}^{l\textrm{,}h}}^{\top}+\bm{M}\right)\bm{V}^{l\textrm{,}h}+\bm{v}^{*l\textrm{,}h}$ 17.53 3.73
@@ 7.3 Attention biases
Learnable biases in attention. In Section 3.1, we have shown that the first token acts as a bias: its key $\bm{k}_{1}^{l\textrm{,}\,h}$ is distributed in a different manifold and its value $\bm{v}_{1}^{l\textrm{,}\,h}$ has small $\ell_{2}$ norm. Xiao et al. (2023b) considered a learnable sink token in each chunk before the input tokens during LM pre-training. As this token is fixed in the first token, this could be considered as implicitly introducing biases $\bm{k}^{*l,h}$, $\bm{v}^{*l,h}$, $\bm{q}^{*l,h}$ in attention, as shown in the second row in Table 4. These biases are the MLP output of $\bm{x}^{*}\bm{W}_{E}$. Sun et al. (2024) directly introducing $\bm{k}^{*l,h}$ and $\bm{v}^{*l,h}$ as learnable parameters in attention (KV biases). They found that this design could alleviate the massive activations. Considering the important role of $\bm{k}^{*l,h}$ and small $\ell_{2}$-norm of $\bm{v}^{*l,h}$ , we propose introducing only the key biases $\bm{k}^{*l,h}$ and fix value biases $\bm{v}^{*l,h}=\bm{0}$ (K biases). As a control, we also consider only adding value biases (V biases). The formulations of all these attention designs are shown in Table 4.
LMs need key biases. After evaluating the LMs with setups in Table 4, we first observe that these LMs have comparable model performance. Moreover, as long as there are key biases $\bm{k}^{*l\textrm{,}\,h}$, attention sink disappears on the first token but on the biases. From the setup of K biases, we reaffirm that the sink token acts as key biases, storing extra attention scores, which could be completely non-informative and not contribute to the value computation. It is worth mentioning that the introduced learnable sink token, KV biases, and V biases become part of model parameters in LMs. Removing them will lead to no attention sink in the first position, but a significant drop in model performance. In Figure 7(Middle), we also find that the LM with a learnable sink token has massive activations in $\bm{x}^{*}$. While LMs with KV biases and K biases have no massive activations.
Table 5: Larger $\ell_{2}$-norm of fixed $\bm{v}^{*l\textrm{,}h}$ results in LMs allocating more attention on $\bm{x}_{1}$ instead of $\bm{k}^{*l\textrm{,}h}$.
@@
Beyond all zeros in V biases. In the setup of K biases, we fix $\bm{v}^{*l,h}=\bm{0}$. We wonder whether the fixed values of $\bm{v}^{*l,h}$ could affect the attention sink. We consider, $\bm{v}^{*l,h}=m\bm{v}^{\prime}$ or $\bm{v}^{*l,h}=m\bm{v}^{\prime\prime}$, where $m$ is the controllable $\ell_{2}$ norm and $\bm{v}^{\prime}=[1,0,0,..,0]$ and $\bm{v}^{\prime\prime}=[1,1,1,..,1]/\sqrt{d_{h}}$. As shown in Table 5, with larger $\ell_{2}$-norm of $\bm{v}^{*l,h}$, attention sink shifts from $\bm{k}^{*l,h}$ to the first token. Intuitively, it is difficult for LMs to remove the effects of $\bm{v}^{*l,h}$ with larger $\ell_{2}$-norm in model predictions. Then they opt to optimize the keys and values of the first token to save extra attention.
Design space of biases. In Table 12(Right), the LM with head-sharing KV biases tends to shift the sink from $\bm{k}^{*l,h}$ back to the first token. While the one with head-sharing K biases is less affected. In Table 13, even with small learnable dimensions for key biases, they can still absorb large attention.
@@ 7.4 Attention operation
General formulation of attention. In the last section, we realize that LMs need key biases to save extra attention. This motivates us to explore whether such a property is related to the dependence among attention scores due to the softmax operation. First, the attention output for the $i$-th token can be generalized as: $\bm{v}_{i}^{\dagger}=\left(\sum_{j^{\prime}=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j^{\prime}}))\right)^{-1}\sum_{j=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j}))\bm{v}_{j}=\bm{Z}_{i}^{-1}\sum_{j=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j}))\bm{v}_{j}$, where we omit the PE, and block/head indexes for simplicity. $\bm{Z}_{i}$ is a normalization term and $\varphi(\cdot)$ is a kernel function. Normally, $\bm{Z}_{i}=\sum_{j^{\prime}=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j}))$. For softmax attention, $\varphi(\cdot)$ is an identity kernel and $\textrm{sim}(\bm{q}_{i}\textrm{,}\,\bm{k}_{j})=\textrm{exp}(\bm{q}_{i}\bm{k}_{j}^{\top}/\sqrt{d_{h}})$.
Normalization. In Table 14(Left) and Figure 29, we show that modifying normalization may result in less obvious attention sink but does not stop its emergence. So we consider removing the normalization. Since the exponential function in softmax tends to explode without normalization, we replace it with sigmoid or elu plus one. When evaluating the attention sink, we compute the proxy attention scores by using the term $\bm{Z}_{i}=\sum_{j^{\prime}=1}^{i}\textrm{sim}(\varphi(\bm{q}_{i})\textrm{,}\,\varphi(\bm{k}_{j^{\prime}}))$ for attention sink metric $\textrm{Sink}_{1}^{\epsilon}$. As shown in Table 6, without normalization, LMs still have comparable validation loss but no attention sink. With normalization, attention sink also emerges in LMs with sigmoid attention.
@@
$\textrm{exp}(\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}})$ $\sum_{j^{\prime}=1}^{i}\textrm{exp}(\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{\top}}{\sqrt{d_{h}}})$ 18.18 3.73
$\textrm{sigmoid}(\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}})$ $1$       0.44 3.70
$\textrm{sigmoid}(\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}})$ $\sum_{j^{\prime}=1}^{i}\textrm{sigmoid}(\frac{\bm{q}_{i}\bm{k}_{j^{\prime}}^{\top}}{\sqrt{d_{h}}})$ 30.24 3.74
$\textrm{elu}(\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}})+1$ $1$       0.80 3.69
$\frac{(\textrm{elu}(\bm{q}_{i})+1)(\textrm{elu}(\bm{k}_{j})+1)^{\top}}{\sqrt{d_{h}}}$ $\sum_{j^{\prime}=1}^{i}\frac{(\textrm{elu}(\bm{q}_{i})+1)(\textrm{elu}(\bm{k}_{j^{\prime}})+1)^{\top}}{\sqrt{d_{h}}}$    53.65 4.19
$\frac{\bm{q}_{i}\bm{k}_{j}^{\top}}{\sqrt{d_{h}}}$ $1$      0.00 3.99
$\frac{\textrm{mlp}(\bm{q}_{i})\textrm{mlp}(\bm{k}_{j})^{\top}}{\sqrt{d_{h}}}$ $\textrm{max}\left(\left|\sum_{j^{\prime}=1}^{i}\frac{\textrm{mlp}(\bm{q}_{i})\textrm{mlp}(\bm{k}_{j^{\prime}})^{\top}}{\sqrt{d_{h}}}\right|\textrm{,}\,1\right)$      0.19 3.85
$\frac{\textrm{mlp}(\bm{q}_{i})\textrm{mlp}(\bm{k}_{j})^{\top}}{\sqrt{d_{h}}}$ $1$      0.74 3.91
@@
Scale up to 1B parameters. We compare model behaviors of 1B LMs with softmax attention and sigmoid attention (without normalization). Specifically, the latter achieves a validation loss of 3.10, slightly larger than the 3.07 achieved by the former. However, the attention sink metric significantly drops from 45.11% to near zero: $\textrm{Sink}_{1}^{\epsilon}=\textrm{2.46\%}$% using the proxy attention scores. Meanwhile, as present in Figure 8(Right), LLMs with sigmoid attention have no massive activations. Furthermore, they have no issues of training stability during continued supervised fine-tuning in Appendix E.